Glass Houses Pty has found that the cost of fitting new sliding doors is normally distributed, with an average cost per door of R12 750, and a standard deviation of R1 095. A sample of 105 new windows is randomly selected. Assuming normal distribution: What is the probability that the average cost of fitting the doors in

the sample is less than R13 000? Interpret your answer.


1
Expert's answer
2021-10-26T07:10:13-0400

Let "\bar{X}=" the average cost of fitting the doors in the sample: "\bar{X}\sim N(\mu_{\bar{X}}, \sigma_{\bar{X}}^2)"

"\mu_{\bar{X}}=\mu, \sigma_{\bar{X}}^2=\sigma^2/n"

Given "\mu=12750, \sigma=1095, n=105."


"P(\bar{X}<13000)=P(Z<\dfrac{13000-12750}{1095/\sqrt{105}})"

"\approx P(Z<2.3395)\approx0.990345"

The probability that the average cost of fitting the doors in the sample is less than R13 000 is "0.990345."